# Enhancing Material Screening at Boulby Underground Laboratory with XIA UltraLo-1800 Alpha Particle Detectors

### v3

@@ Abstract
The Boulby UnderGround Screening (BUGS) facility, located at the Boulby Underground Laboratory, has significantly advanced its material screening capabilities by installing two XIA UltraLo-1800 alpha particle detectors. This study presents a comprehensive evaluation of one of these detectors, operated 1,100 meters underground at the Boulby Underground Laboratory, which provides significant shielding from cosmic radiation and maintains a low ambient radon activity of 2.30 $\pm$ 0.03 Bq/m3. Our evaluation focuses on energy reconstruction accuracy, background radiation rates, and operational stability. The XIA UltraLo-1800 detector demonstrates remarkable stability in energy reconstruction, with less than 0.1 MeV variation over four years. Moreover, the implementation of a graphite-filled PTFE liner in the sample tray resulted in a significant reduction in background radiation levels compared to measurements with the original stainless steel tray, achieving an average activity of 0.15 $\pm$ 0.01 $\alpha$/cm2/khr. Copper sample assays, performed before and after radon exposure, demonstrated the detector’s ability to accurately identify and quantify 210Po contamination. By implementing the robust cleanliness procedures and protocols described in this article, we observed a reduction in 210Po activity from 0.504 $\pm$ 0.022 mBq to 0.336 $\pm$ 0.013 mBq, highlighting the crucial role of refined cleaning methods in minimizing background for sensitive experiments. Additionally, observations of elevated background activity levels post-high-activity sample measurements illustrate the need for careful management of assay conditions and environment to maintain low background levels. These results highlight the potential of the XIA UltraLo-1800 in enhancing the precision of material assays essential for reducing background interference in rare event experiments.
@@ 2.1 Detector Calibration and Energy Reconstruction
The Boulby UltraLo-1800 was routinely calibrated using a custom-made 210Po alpha-particle source provided by Spectrum Techniques [20]. 210Po decays by emitting a 5.304 MeV alpha particle with a branching ratio of essentially 100% [21]. To better understand the detector’s response to various materials and surface contamination, we constructed comprehensive simulations using the GEANT4-v10.5.1 toolkit [22]; G4EmStandardPhysics, PhysListEmPenelope, and PhysListEmLivermore were used to model electromagnetic processes, while the G4HadronElasticPhysicsHP was used to model hadronic interactions [23, 24]. The geometry of the XIA UltraLo-1800 detector was implemented in the simulation, including all physical aspects, such as the active volume, anode, and guard rail, as well as its immediate surroundings. The 210Po source was modelled as an isotropic source positioned at the centre of the sample tray.
@@ 2.2 Background Measurements
The detector used for this study has been fully operational since January 2018. The sample tray is a critical component of the detector, as it holds the materials being assayed and determines the sensitive area exposed to the detector. To understand the inherent background measurements, we started by conducting measurements with the manufacturer-supplied stainless steel sample tray empty. A photograph of the stainless steel sample tray is shown in the left-hand panel of Figure 3. Several measurements were made, shown as the filled red data points in Figure 4, indicating an average activity of 1.28 $\pm$ 0.03 $\alpha$/cm2/khr. This result is in good agreement with results from other UltraLo-1800 facilities [16, 17, 18].
@@
Figure 4: Long-term stability study of the UltraLo-1800 event rate with different tray configurations. The filled red data points show the initial operation with the stainless steel sample tray. The filled blue data points showed a reduction in event rate when a graphite-filled PTFE liner was inserted into the tray. Unfilled blue squares represent measurements made immediately following sample assays, indicating the necessity of a purge period to return to the low background configuration. The study affirms the operational stability of the UltraLo-1800 and the effectiveness of the PTFE liner in reducing background interference.
@@
Our long-term study of the UltraLo-1800 with this PTFE liner inserted (but otherwise empty) is presented as the filled blue squares in Figure 4. The mean emissivity is measured to be 0.14 $\pm$ 0.01 $\alpha$/cm2/khr; individual measurements have a standard deviation of 0.01, indicating the stability of these measurements.
It is also shown in Figure 4 as the unfilled blue squares are measurements made with an identical detector configuration performed immediately following runs in which a sample had been assayed. As we can see, following a sample measurement, a significant purge period (typically 48 hours) is necessary to return the instrument to its lowest background configuration.
@@
Further efforts to reduce background were made by replacing the tray with electroformed copper from the Pacific Northwest National Laboratory (PNNL). For transport from PNNL to Boulby, this copper sheet had been placed within three layers of heat-sealed 50-micrometre thick metallized mylar bags back-filled with nitrogen at atmospheric pressure. This packaging prevents exposure to radon and thus build up of surface plate-out. At Boulby, the electroformed copper was placed at the centre of the UltraLo-1800 and measured in ‘wafer mode’, in which the detector’s sensitive area is limited to that of the sample. An emissivity of 0.13 $\pm$ 0.04 $\alpha$/cm2/khr was recorded.
@@ 3 Surface and Bulk Alpha Measurements
Figure 6: Spectra illustrating simulations of energy depositions from surface and bulk 210Po alpha particles in copper. (Left) Detected alpha particle energy spectra from simulations of 105 210Po decays in successive 1 $\mu$m thick slices of copper. (Right) Simulated combined surface and bulk 210Po alpha particle energy spectrum in copper.
@@ 3.1 Assay of Copper Plate-out and Bulk Activity
Name Description 210Po Activity ($\mu$Bq/cm2) Implications

### v4

@@ Abstract
The Boulby UnderGround Screening (BUGS) facility, located at the Boulby Underground Laboratory, has significantly advanced its material screening capabilities by installing two XIA UltraLo-1800 alpha particle detectors. This study presents a comprehensive evaluation of one of these detectors, operated 1,100 meters underground at the Boulby Underground Laboratory, which provides significant shielding from cosmic radiation and maintains a low ambient radon activity of 2.30 $\pm$ 0.03 Bq/m3. Our evaluation focuses on energy reconstruction accuracy, background radiation rates, and operational stability. The XIA UltraLo-1800 detector demonstrates remarkable stability in energy reconstruction, with less than 0.1 MeV variation over four years. Moreover, the implementation of a graphite-filled PTFE liner in the sample tray resulted in a significant reduction in background radiation levels compared to measurements with the original stainless steel tray, achieving an average activity of 0.15 $\pm$ 0.01 $\alpha$/cm2/khr (kilo-hour, or 1000 hours). Copper sample assays, performed before and after radon exposure, demonstrated the detector’s ability to accurately identify and quantify 210Po contamination. By implementing the robust cleanliness procedures and protocols described in this article, we observed a reduction in 210Po activity from 0.504 $\pm$ 0.022 mBq to 0.336 $\pm$ 0.013 mBq, highlighting the crucial role of refined cleaning methods in minimizing background for sensitive experiments. Additionally, observations of elevated background activity levels post-high-activity sample measurements illustrate the need for careful management of assay conditions and environment to maintain low background levels. These results highlight the potential of the XIA UltraLo-1800 in enhancing the precision of material assays essential for reducing background interference in rare event experiments.
@@ 2.1 Detector Calibration and Energy Reconstruction
The Boulby UltraLo-1800 was routinely calibrated using a custom-made 210Po alpha-particle source, 1 inch in diameter, provided by Spectrum Techniques [21]. 210Po decays by emitting a 5.304 MeV alpha particle with a branching ratio of essentially 100% [22]. To better understand the detector’s response to various materials and surface contamination, we constructed comprehensive simulations using the GEANT4-v10.5.1 toolkit [23]; G4EmStandardPhysics, PhysListEmPenelope, and PhysListEmLivermore were used to model electromagnetic processes, while the G4HadronElasticPhysicsHP was used to model hadronic interactions [24, 25]. The geometry of the XIA UltraLo-1800 detector was implemented in the simulation, including all physical aspects, such as the active volume, anode, and guard rail, as well as its immediate surroundings. The 210Po source was modelled as an isotropic source positioned at the centre of the sample tray.
@@ 2.2 Background Measurements
The detector used for this study has been fully operational since January 2018. The sample tray is a critical component of the detector, as it holds the materials being assayed and determines the sensitive area exposed to the detector. To understand the inherent background measurements, we started by conducting measurements with the manufacturer-supplied stainless steel sample tray empty. A photograph of the stainless steel sample tray is shown in the left-hand panel of Figure 3. Several measurements were made, shown as the filled red data points in Figure 4, indicating an average rate of 1.28 $\pm$ 0.03 $\alpha$/cm2/khr. This result is in good agreement with results from other UltraLo-1800 facilities [17, 18, 19].
@@
Figure 4: Long-term stability study of the UltraLo-1800 event rate with different tray configurations. The filled green data points show the initial operation with the stainless steel sample tray. The filled yellow squares represent background measurements taken after a 48-hour purge; early data points (before the start of 2019) may reflect the system stabilizing after opening. The unfilled yellow squares represent measurements made immediately following sample assays, indicating the necessity of a purge period to return to the low background configuration. The study affirms the operational stability of the UltraLo-1800 and the effectiveness of the PTFE liner in reducing background interference.
@@
Our long-term study of the UltraLo-1800 with this PTFE liner inserted (but otherwise empty) is presented as the filled yellow squares in Figure 4. The mean emissivity is measured to be 0.14 $\pm$ 0.01 $\alpha$/cm2/khr; individual measurements have a standard deviation of 0.01, indicating the stability of these measurements. It should be noted that some early background measurements, particularly those taken at the beginning of 2018, exhibit higher values than those observed later in the study. This is likely due to the system stabilizing after being opened to the atmosphere.
It is also shown in Figure 4 as the unfilled yellow squares are measurements made with an identical detector configuration performed immediately following runs in which a sample had been assayed. As we can see, following a sample measurement, a significant purge period (typically 48 hours) is necessary to return the instrument to its lowest background configuration.
@@
Further efforts to reduce background were made by replacing the tray with electroformed copper from the Pacific Northwest National Laboratory (PNNL). For transport from PNNL to Boulby, this copper sheet had been placed within three layers of heat-sealed 50 $\mu$m thick metallized mylar bags back-filled with nitrogen at atmospheric pressure. This packaging prevents exposure to radon and thus build up of surface plate-out. At Boulby, the electroformed copper was placed at the centre of the UltraLo-1800 and measured in ‘wafer mode’, in which the detector’s sensitive area is limited to that of the sample. An emissivity of 0.13 $\pm$ 0.04 $\alpha$/cm2/khr was recorded.
@@ 3 Surface and Bulk Alpha Measurements
Figure 6: Spectra illustrating simulations of energy depositions from surface and bulk 210Po alpha particles in copper. (Top) Detected alpha particle energy spectra from simulations of 105 210Po decays in successive 1 $\mu$m thick slices of copper. (Bottom) Simulated combined surface and bulk 210Po alpha particle energy spectrum in copper.
@@ 3.1 Assay of Copper Plate-out and Bulk Activity
Name Description 210Po Surface Activity ($\mu$Bq/cm2) Implications